OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1).

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

@MoonlitFate Hey, girly. Do you mind helping me with this quesion?

OpenStudy (moonlitfate):

Okay, so to figure out what equations would be parallel or perpendicular to another depends on their slope. Perpendicular lines have slopes that opposite reciprocals of each other. Parallel lines are going to have the same slope.

OpenStudy (moonlitfate):

So the slope y = x +5 is 1. It's in slope-intercept form: y = mx+b, whatever is next to x is the slope. The leading coefficient of x is just 1. :) A parallel line to y = x+5 would be the negative reciprocal of 1. You can remember that by just flipping the slope and changing its sign. So, the negative reciprocal of 1 would be -1. (1 = 1/1 -> so flip it and switch its sign).

OpenStudy (anonymous):

So it will be B?

OpenStudy (moonlitfate):

It would be C; sorry for the long time to reply. Took me a bit to figure out since it's been a while since I've done Algebra I things, :p

OpenStudy (moonlitfate):

Basically to find that out-- since you know what the slopes have to be for parallel and perpendicular lines, you just have to use slope-intercept form: y=mx+b to find what b would be in those equations. b is the y-intercept. Then you just plug the slope and y-intercept back in.

OpenStudy (moonlitfate):

Line that's perpendicular to y = x +5 and passes through (-2, -1). m = -1 To find the y-intercept use the point (-2, -1) and slope-intercept formula: y = mx +b ; -1=-1(-2) =b -1 = 2+b -2-1 = b -3 = b To find the equation of the line that's perpendicular, plug in the slope and y-intercept back in to slope-intercept formula y = -x -3